# Experiential Semantic Information and Brain Alignment: Are Multimodal Models Better than Language Models?

### v1

@@ 6 Discussion and Conclusions
Another noteworthy result is that the correlations with fMRI responses we found for SimCSE, MCSE and BERT are higher than those achieved by the computational models (GloVe and Word2vec) tested in Fernandino et al. (2022) (see Appendix C.3). This finding aligns with previous work showing that transformer-based architectures are more predictive of brain responses during language processing than word-level embedding models and recurrent neural networks Schrimpf et al. (2021).
Finally, one last interesting finding was that the LMs and, to a larger extent, the VLMs we tested learn brain-relevant semantic information beyond that captured by Exp48. This partially echoes the results by Carota et al. (2024), with the difference that the computational model included in their study was strictly distributional.
@@
Our study has several implications for future work. First, it invites caution against assuming that technical innovations allowing models to solve additional downstream tasks should necessarily make them more ‘human-like’. Second, it indicates that there is significant room for improving current computational language models so that they learn the brain-relevant experiential information they currently lack—how to concretely achieve this remains an open question.
@@
Appendix A Limitations
Despite our efforts to develop a robust experimental design, a few limitations should still be considered when drawing conclusions from our results. First, the experiential model we used—while showed to reflect brain-relevant information—relies on pre-defined dimensions that are, to some extent, abstract and high-level. In this sense, it may fail to capture low-level perceptual information relevant for modelling human word representations and, perhaps, learnt by computational models.
Another limitation is that our study focuses on single words as opposed to longer text passages. To some extent, our experiments suggest that this may affect machine language processing, as we found representations by computational models to be more aligned with brain responses when words are passed within sentence templates vs. in isolation (see also Appendix C.1). Similarly, the isolated-word setup is likely to play a role human language processing—as insightfully observed by Zwaan (2014), context is crucial in determining the mental simulations people engage during language comprehension and the amount of perceptual detail they contain.
Lastly, two additional limitations are related to the models included in our study. The first is that, while SimCSE, MCSE and CLAP are comparable in their learning objectives and architecture, they still differ in the amount of fine-tuning data. While, in principle, this could be problematic, we observed that SimCSE—despite being fine-tuned on less sentences than MCSE—still proves to be more Exp48- and brain-aligned. As for CLAP, we acknowledge that the smaller amount of fine-tuning audio-caption pairs, together with the fact that it did not optimise a SimCSE-like objective jointly with the CLIP-like one, may have played a role in its poor performance.
The second limitation related to our model selection is that we only considered architectures fine-tuned contrastively, which are known to reflect high-level correspondences between objects present in images and captions, while failing to capture more fine-grained relations between modalities Hendricks and Nematzadeh (2021); Parcalabescu et al. (2022); Thrush et al. (2022); Liu et al. (2023); Chen et al. (2023); Bavaresco et al. (2024). Therefore, it could be that different trends would emerge if repeating the experiments with computational models trained with different objectives.
Appendix B Sentence Templates
@@
Appendix C Additional RSA Results
@@
C.1 Single-word vs. contextualised representations
@@
C.2 Layer-wise RSA results
@@
C.3 RSA with additional baselines

### v2

@@ 6 Discussion
An interesting result was that the correlations with fMRI responses we observed for SimCSE, MCSE and BERT are higher than those achieved by the computational models (GloVe and Word2vec) tested by Fernandino et al. 2022 (see Appendix B.3). This finding aligns with previous work showing that transformer-based architectures are more predictive of brain responses during language processing than word-level embedding models and recurrent neural networks Schrimpf et al. (2021). In addition, we found that the LMs and, to a larger extent, the VLMs we tested learn brain-relevant semantic information beyond that captured by Exp48. This partially echoes the results by Carota et al. (2024), with the difference that the computational model included in their study was strictly distributional.
Information captured by Exp48
While the ability of Exp48 to model brain responses has been validated by previous research, it may still be a suboptimal model of perceptual experience for two main reasons. First, all dimensions in Exp48, including the more perceptual ones like Colour or Sound, are somewhat abstract; in this sense, they may fail to capture low-level perceptual information relevant for modelling human word representations and, perhaps, learnt by multimodal models. Second, Exp48 encodes experiential dimensions, e.g., Pleasant or Time, which are not strictly perceptual and may be hard, if not impossible, to learn for vision-language and audio-language models.
Type of stimuli
Our study focuses on single words that are not included in longer text passages. To some extent, our results suggest that this may affect machine language processing; indeed, we found that embedding words in sentences, as opposed to passing them to the models as is, yields more brain-aligned representations (see also Appendix B.1). In a similar vein, the amount of context may influence human language processing: As observed by Zwaan (2014), context determines the perceptual detail of the mental simulations people engage during language comprehension. Therefore, it may be that the nouns used in the fMRI experiment did not prompt multimodal semantic knowledge enough for it to be detected in our study.
7 Conclusions
Our study provides an in-depth comparison between multimodal and language-only architectures in their ability to capture experiential semantic information and alignment with brain responses. Contrary to common assumptions, we found multimodal models to produce word representations less brain-aligned and experience-informed than language-only models.
@@ 7 Conclusions
These results have several implications for future work. First, they invite caution against assuming that technical innovations allowing models to solve additional downstream tasks should necessarily make them more ‘human-like’. Second, they indicate that there is significant room for improving current computational language models so that they learn the brain-relevant experiential information they currently lack—how to concretely achieve this remains an open question.
Limitations
Our experimental setup focuses exclusively on contrastive models which are not state-of-the-art for both linguistic and multimodal downstream tasks. More recent architectures pretrained autoregressively—e.g., models from the LLaVA family Liu et al. (2024), Molmo Deitke et al. (2024), or Qwen2.5-VLBai et al. (2025)—may exhibit different patterns. However, the complexity of their pre-training and fine-tuning steps makes it hard to set up a controlled comparison ruling out factors such as the amount of training data or training objectives. We therefore explicitly decided to not include this type of model in our investigation. This decision was further informed by preliminary evidence that generative vision-language models achieving stronger performance on downstream tasks are less brain-aligned than previous architectures Bavaresco et al. (2024).
Acknowledgments
We thank the members of the Dialogue Modelling Group (DMG) from the University of Amsterdam and Lorenzo Proietti for the helpful feedback provided at different stages of this project.
This project was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 819455).
@@ Acknowledgments
Appendix A Sentence Templates
@@
Appendix B Additional RSA Results
@@
B.1 Single-word vs. contextualised representations
@@
B.2 Layer-wise RSA results
@@
B.3 RSA with additional baselines